# Is Complex Query Answering Really Complex?

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
over a KG such as FreeBase (Bollacker et al., 2008), one would need to first intersect the set of movies found on Blue Ray and the ones shot in New York City, and then link these intermediate candidate answers to another entity, i.e., an actor participating in it. However, the answers computed in this way may not include entities that are unreachable if missing links are present. To deal with the unavoidable incompleteness of real-world KGs, ML methods were developed to solve CQA in the presence of missing links. The current state-of-the-art (SoTA) for CQA are neural query answering models,(Arakelyan et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023; Ren et al., 2023; Galkin et al., 2024b), mapping queries and KGs (i.e., entities and relation names) into a unified latent space. Performance measured on de-facto-standard benchmarks such as $\mathsf{FB15k237}$ (Toutanova & Chen, 2015) and $\mathsf{NELL995}$ (Xiong et al., 2017) suggests impressive progress achieved in recent years on CQA on queries having different structures, and hence posing apparently different levels of difficulty to be answered.
@@
The difficulty of a benchmark relates to the size and structural complexity [^1] Not to be confused with the computational complexity of query answering (Dalvi & Suciu, 2012). of its queries, and several query “types” have been proposed (Ren et al., 2020), each involving a different combination of logical operators—conjunctions, disjunctions, and requiring to traverse a number of missing links that generally increases with the number of logical conditions imposed (Figures 1 and 2). For instance, the query $q_{1}$ is an example of a “2i1p” query type, since it comprises an intersection of two entity sets (2i) followed by a path [^2] Also known as projection in related works (Ren et al., 2023). of length one (1p). In this paper, we argue that the perception of progress on CQA benchmarks has been distorted by implicit assumptions in these benchmarks. We start by noting how both in $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, the vast majority of queries of a complex type simplify to one simpler type, thanks to the fact that to answer them, one can leverage links already appearing in the training data. Figure 1 shows an example of how answers to the query $q_{1}$ on $\mathsf{FB15k237}$– which should require predicting three missing links – require the same effort associated with simpler types involving fewer links. In fact, in $\mathsf{FB15k237}$, the answer “K. Dunst” can be retrieved by predicting just one missing link when leveraging the training links. In this sense, a 2i1p query reduces to a “1p” query, i.e., the much simpler task of link prediction (see Sec. 2). Similarly, the answers “J. Bryant” and “K. Chandler” can be retrieved by predicting two links instead of three. Only the answer “A. Serkis” requires predicting three missing links, and thus follows the intuitive “hardness” associated to the type 2i1p, i.e., it is the hardest when no information is available. As we note here, there is a spectrum of hardness, predicting “K. Dunst” will be easier than answering “K.Chandler”, which in turn is easier than saying “A. Serkis”. However, current CQA benchmarks greatly overrepresent the easy query answers over the others – the “K. Dunst”-like answers are the vast majority – and this inflates performance as SoTA models are prone to memorize training links. In this paper, we analyze this issue and propose a new set of benchmarks that are balanced and cover the full spectrum of hardness in CQA.
@@ 2 KGs and Complex Query Answering
Different queries are categorized into different types based on the structure of their corresponding logical sentence (Xiong et al., 2017). This taxonomy should imply that queries of the same type share the same “hardness”, i.e., the level of difficulty to be answered, and different types correspond to tasks that map to more or less complex reasoning tasks. As we will show in the next sections, this is not the case, and to fully grasp the actual hardness to answer a query one needs to account for the “leakage” of intermediate links. We now review the most common query types used in CQA. The simplest CQA task is link prediction, i.e., answering a query of the form:
@@
Note that despite their dissimilar syntaxes and the associated sub-graphs (Figure 2), the query type 2u should be as difficult as 1p, as to answer the first it suffices to match a single link correctly. Similarly, 2u1p is as complex as 2p. The fact that 2u and 2u1p are reported to be harder to solve in practice than 1p and 2p (Ren et al., 2020) is due to the way standard benchmarks are built, which we discuss in Sec. 4, and the way in which CQA is evaluated, discussed next. App. B details further query types, including negation.
Standard evaluation. Given a KG $G$ and a query $q$, answering $q$ boils down to a graph matching problem (Hogan, 2020) if we assume that all the meaningful links are already in $G$. Instead, if $G$ is incomplete, we will need to predict missing links while answering $q$. Many ML approaches to CQA, reviewed in Sec. 5, therefore assume a distribution over possible links (Loconte et al., 2023), requiring probabilistic reasoning. To evaluate them, standard benchmarks such as $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ artificially divide $G$ into $G_{\operatorname{train}}$ and $G_{\operatorname{test}}$, treating the triples in the latter as missing links. This splitting process is done uniformly at random, a procedural choice that can alter the measured performance, as we discuss next. CQA is generally treated as a ranking task, counting how many true candidate answers to a query are ranked higher than non-answer ones. Let the rank of a query answer (QA) pair $(q,t)$ be $\operatorname{rank}(q,t)$, the performance for each query type is calculated as the mean reciprocal rank (MRR), i.e.:
@@ 3 What is the real “Hardness” of CQA in incomplete KGs?
Given a query $q$ and every answer set of candidate answers $t\in\mathcal{E}_{q}$, we define the reasoning tree of each QA pair $(q,t)$, as the directed acyclic graph starting from the anchor entities of $q$ to the target entity $t$, whose relational structure matches the query graph. Figure 1 provides examples of different reasoning trees for four different answers to the same query. There, we highlight whether a link belongs to $G_{\text{train}}$ or not, i.e., it is a missing link. We assess the hardness of each answer $t\in\mathcal{E}_{q}$, by analyzing the composition of the reasoning tree required to predict $t$. As to answering a $(q,t)$ pair, multiple reasoning trees are possible, so we select the one with the smallest number of missing links and the fewest number of hops. A $(q,t)$ pair is trivial if the answer $t$ can be entirely retrieved from $G_{\text{train}}$, i.e., there exists at least one reasoning tree where all the links in it are seen during training. This type of answer does not need probabilistic inference and hence is filtered out from current CQA benchmarks (Ren & Leskovec, 2020), which only consider non-trivial $(q,t)$ pairs, which we call inference-based pairs. However, inference-based pairs do not need all links in their reasoning tree to be predicted as, by definition, it is sufficient that at least one link in the tree is missing. Therefore, we define a $(q,t)$ pair to be a partial-inference pair if its reasoning tree contains at least one link in $G_{\text{train}}$ and at least one link present in $G_{\text{test}}$. Alternatively, a QA pair whose reasoning tree contains only links present in $G_{\text{test}}$ is called a full-inference pair. We claim that inference-based queries encompass a spectrum of hardness, with full-inference queries being the hardest and queries that have only one truly missing link (1p) the easiest. However, current benchmarks flatten this complexity and highlight only performance on the easiest.
@@ 4 How Many “Easy” Queries in Current CQA Benchmarks?
In this section, we systematically analyze the practical hardness of queries from the current CQA benchmarks and answer the following research question: (RQ1) What is the proportion of QA pairs that can be classified as full-inference rather than partial-inference and how easy are the latter? To this end, we consider the CQA benchmarks generated from $\mathsf{FB15k237}$ (based on Freebase) and $\mathsf{NELL995}$ (based on NELL systems (Carlson et al., 2010)) as they are the most used to evaluate SoTA methods for CQA (Ren et al., 2020; Ren & Leskovec, 2020; Arakelyan et al., 2021; Zhang et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023). [^4] We do not consider $\mathsf{FB15k}$ (Bordes et al., 2013) as it suffers from data leakage (Toutanova & Chen, 2015).
Complex queries can be reduced to much simpler types. We group the testing QA pairs $(q,t)$ into query types (Sec. 2), and we further split them based on whether they can be reduced to simpler types after observing the training links in $G_{\operatorname{train}}$ (Sec. 3). Table 1 shows that for both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ the vast majority of $(q,t)$ pairs can be reduced to simpler types. For $\mathsf{FB15k237}$, 86.8% to 98.3% of $(q,t)$ pairs can be reduced to 1p queries, while only 0.1% to 4% require full inference. Similarly, for $\mathsf{NELL995}$, 49.5% to 98.5% of $(q,t)$ pairs map to 1p queries, and only 0.1% to 6% to full inference. For instance, 96.7% of 2i1p $(q,t)$ pairs in $\mathsf{FB15k237}$ can be reduced to 1p queries, 1.8% to 2p, and 1.4% to 2i. However, only 0.1% of these pairs cannot be reduced to any other $(q^{\prime},t)$ pair, i.e., they require full inference in order to be predicted. The only exceptions to this trend are $(q,t)$ pairs where $q$ has a 1p or 2u structure which, by definition, only require the prediction of a single link and therefore cannot be reduced by any other query type. We conduct the same analysis for queries that include negation, considering only the positive reasoning tree (see Sec. A.1 for details). Table A.1 presents results analogous to those in Table 1, showing that the vast majority of $(q,t)$ pairs including negation have training links in their reasoning tree. For example, 95.4% of 3in $(q,t)$ pairs in $\mathsf{FB15k237}$ have training links in their reasoning tree.
@@
Non-existing links for union queries. Finally, we discovered that there are non-existing links, i.e., links that are not in the original KG $G$ and hence neither in $G_{\operatorname{train}}$ or $G_{\operatorname{test}}$, in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ in the reasoning trees of queries involving unions, i.e., the 2u and 2u1p types. Figure A.1 shows some examples. These links violate our definitions of inference ($q$, $t$) pairs, hence, we filter them out, and report only filtered $(q,t)$ pairs in Table 1. More crucially, these non-existing links can alter the performance of solvers for 2u and 2u1p types, as we discuss next.
Takeaway 1. We discourage using current CQA benchmarks as they essentially reflect only the performance of predicting answers where only one link is truly missing (for both hybrid and neural solvers).
@@
Table 2: SoTA performance drops significantly on full-inference QA pairs as shown for the MRR for 3p and 2i1p queries on $\mathsf{FB15k237}$. Compare the MRR achieved on all available testing queries (“ovr”) with that achieved on the 1p reductions (i.e., when only a single link is truly missing): they are extremely close.
@@ 5 What do SoTA Models Perceive as Hard?
Neural CQA methods. Over the years, a significant number of neural models have been proposed for solving the CQA task, see App. C for an overview. Among purely-neural models, we consider four representative approaches that significantly differ in their methodological designs and yield SoTA results compared to other models in their class: (1) Cone Embeddings (ConE; Zhang et al., 2021) is a geometry-based complex query answering model where logical conjunctions and disjunctions are represented as intersections and union of cones. (2) Graph Neural Network Query Executor (GNN-QE; Zhu et al., 2022) decomposes a complex first-order logic query into projections over fuzzy sets. (3) Continuous Query Decomposition (CQD; Arakelyan et al., 2021, 2023), reduces the CQA task to the problem of finding the most likely variable assignment, where the likelihood of each link (1p query) is assessed by a neural link predictor, and logical connectives are relaxed via fuzzy logic operators. (4) UltraQuery (ULTRAQ; Galkin et al., 2024b) is a foundation model for CQA inspired by Galkin et al. (2024a) where links and logical operations are represented by vocabulary-independent functions which can generalize to new entities and relation types in any KG.
Hybrid solvers. So far, we have assessed that current benchmarks are skewed towards easier query types (Table 1) and thus the perceived progress of current SoTA CQA methods boils down to their performance on 1p queries (Table 2). To have an undistorted view of this progress, in the next section, we will devise a benchmark that rebalances all partial-inference as well as full-inference queries types. We argue that in a real-world scenario one has to perform reasoning over both existing links and in the presence of missing ones, hence discarding the information in $G_{\operatorname{train}}$ is wasteful. This, in addition to the fact that purely-neural SoTA CQA methods might implicitly exploit this aspect at test time if they are able to memorize well the entire training KG, pushes us to evaluate also hybrid solvers.
Table 2 presents a portion of the results in terms of MRR for the testing $(q,t)$ pairs in the existing benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, along with the different $(q^{\prime},t)$ pairs they can be reduced to when observing the training KG, grouped by each query type. There, the performance of all SoTA methods consistently drops for query answer pairs that have a higher number of missing links in their reasoning tree. Furthermore, Tables A.5 and A.6 compare the MRR computed on all the available queries originally used in the benchmarks (column “all”), and the scores on those query types that they can be reduced to. There is a high similarity in MRR between the columns “all” and “1p” that is evidence that the good performances of CQA methods are explained by their link prediction performances, as a very high percentage of queries in fact are reduced to 1p queries (see Table 1). A similar conclusion can be drawn for queries including negation, as shown in Table A.2, where the MRR results on partial inference $(q,t)$ pairs are very similar to the overall, and the performance of all SoTA methods consistently drops on full-inference $(q,t)$ pairs.
Note that this happens consistently also for hybrid solvers such as QTO and our CQD-Hybrid: they have a harder time solving QA pairs with more than one missing link. While their performance was consistently surpassing purely-neural solvers, and CQD-Hybrid consistently boosted the performance of CQD, their poor performance on harder QA pairs is not reflected in the overall average score that is usually reported on papers, see Table 2. Moreover, in Table A.3 we highlight that our baseline CQD-hybrid, is able to surpass existing non-hybrid SoTA, which is evidence that a pre-trained neural link predition plus memorization of training links result in SoTA performances for $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. The complete stratified comparison for all solvers is reported in the appendix in Tables A.5 and A.6. However, in a few cases it happens that full-inference results are higher than some partial-inference; in such cases, the reason is to be found in the fact that full-inference $(q,t)$ pairs are very scarce (see Table 1). [^5] We rule out that this is due to the influence of the number of existing entities bounding to existential quantified variables. For example we report such an analysis for CQD in Figures A.3 and D.2. Moreover, we found that no anchor entities nor relation names is predominant in both benchmarks, as shown in Table A.7. This motivates us to create a fully-balanced benchmark in the next section.
@@
Non-existing links for union queries give a false sense of hardness. We remark that for union queries, we filter query answer pairs, as discussed in Sec. 4, removing those with non-existing links in their reasoning trees. In Table 3, we show that if we do not do that, and consider also the pairs with non-existing links (denoted as “2u all”), MRR performance greatly drops. In this way, we reproduce the low performance for 2u queries that the original SoTA baselines reported in their papers. However, as Table 3 shows, this is just an artifact due to including non-filtered pairs while computing Equation 1. With our filtered pairs, instead, we recover the performance of 1p queries, as expected (Sec. 2). Similar conclusions can be drawn for other union queries as reported in Table A.4.
@@
Takeaway 2. $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess the capability of CQA methods to answer complex queries, as most of their QA pairs can be predicted by just predicting a single missing links. This results in highly inflated performance that distorts the perception of progress in the field and pushes the community to improve the performance only on the easiest QA pairs.

### v3

@@ Abstract
complex query answering, knowledge graphs, multi-hop reasoning, neuro-symbolic
@@ 1 Introduction
over a KG such as FreeBase (Bollacker et al., 2008), one would need to first intersect the set of movies found on Blue Ray and the ones shot in New York City, and then link these intermediate candidate answers to another entity, i.e., an actor participating in it. However, if some links are missing in the KG, not all relevant actor entities can be immediately reachable. To deal with the unavoidable incompleteness of real-world KGs, ML methods were developed to solve CQA in the presence of missing links. The current state-of-the-art (SoTA) for CQA are neural query answering models (Arakelyan et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023; Ren et al., 2023; Galkin et al., 2024b), mapping queries and KGs (i.e., entities and relation names) into a unified latent space. Performance measured on de-facto-standard benchmarks such as $\mathsf{FB15k237}$ (Toutanova & Chen, 2015) and $\mathsf{NELL995}$ (Xiong et al., 2017) suggests impressive progress achieved in recent years on CQA on queries having different structures, and hence posing apparently different levels of difficulty to be answered.
@@
The difficulty of a benchmark relates to the size and structural complexity [^1] Not to be confused with the computational complexity of query answering (Dalvi & Suciu, 2012). of its queries, and several query “types” have been proposed (Ren et al., 2020), each involving a different combination of logical operators—conjunctions, disjunctions, and requiring to traverse a number of missing links that generally increases with the number of logical conditions imposed (Figures 1 and 2). For instance, the query $q_{1}$ is an example of a “2i1p” query type, since it comprises an intersection of two entity sets (2i) followed by a path [^2] Also known as projection in related works (Ren et al., 2023). of length one (1p). In this paper, we argue that the perception of progress on CQA benchmarks has been distorted by implicit assumptions in these benchmarks. We start by noting how both in $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, the vast majority of queries of a complex type simplify to one simpler type, thanks to the fact that to answer them, one can leverage links already appearing in the training data. Figure 1 shows an example of how answers to the query $q_{1}$ on $\mathsf{FB15k237}$– which should require predicting three missing links – require the same effort associated with simpler types involving fewer links. In fact, in $\mathsf{FB15k237}$, the answer “K. Dunst” can be retrieved by predicting just one missing link when leveraging the training links. In this sense, a 2i1p query reduces to a “1p” query, i.e., the much simpler task of link prediction (see Sec. 2). Similarly, the answers “J. Bryant” and “K. Chandler” can be retrieved by predicting two links instead of three. Only the answer “A. Serkis” requires predicting three missing links, and thus follows the intuitive “hardness” associated to the type 2i1p, i.e., it is the hardest when no information is available. As we note here, there is a spectrum of hardness, predicting “K. Dunst” will be easier than answering “K.Chandler”, which in turn is easier than saying “A. Serkis”. However, current CQA benchmarks greatly overrepresent the easy query answers over the others – the “K. Dunst”-like answers are the vast majority – and this inflates performance as SoTA models are known to be prone to memorize training links. In this paper, we analyze this issue and propose a new set of benchmarks that cover the full spectrum of hardness in CQA.
@@ 2 KGs and Complex Query Answering
Different queries are categorized into different types based on the structure of their corresponding logical sentence (Xiong et al., 2017). This taxonomy should imply that queries of the same type share the same “hardness”, i.e., the level of difficulty to be answered, and different types correspond to tasks that map to more or less complex reasoning tasks. As we will show in the next sections, this is not the case, and to fully grasp the actual hardness to answer a query one needs to account for the “leakage” of intermediate training links. We now review the most common query types used in CQA. The simplest CQA task is link prediction, i.e., answering a query of the form:
@@
Note that despite their dissimilar syntaxes and the associated sub-graphs (Figure 2), the query type 2u should be as difficult as 1p, as to answer the first it suffices to match a single link correctly. Similarly, 2u1p should be as complex as 2p. The fact that 2u and 2u1p are reported to be harder to solve in practice than 1p and 2p (Ren et al., 2020) is due to the way standard benchmarks are built, which we discuss in Sec. 4, and the way in which CQA is evaluated, discussed next. App. B details further query types, including negation.
Standard evaluation. Given a KG $G$ and a query $q$, answering $q$ boils down to a graph matching problem (Hogan, 2020) if we assume that all the meaningful links are already in $G$. Instead, if $G$ is incomplete, we will need to predict missing links while answering $q$. Many ML approaches to CQA, reviewed in Sec. 5, therefore assume a distribution over possible links (Loconte et al., 2023), thus requiring some form of probabilistic reasoning. To evaluate them, standard benchmarks such as $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ artificially divide $G$ into $G_{\operatorname{train}}$ and $G_{\operatorname{test}}$, treating the triples in the latter as missing links. This splitting process is done uniformly at random, a procedural choice that is inherited from link prediction benchmarks (Ruffinelli et al., 2020) and that can alter the measured performance, as we discuss next. CQA is generally treated as a ranking task, counting how many true candidate answers to a query are ranked higher than non-answer ones. Let the rank of a query answer (QA) pair $(q,t)$ be $\operatorname{rank}(q,t)$, the performance for each query type is calculated as the mean reciprocal rank (MRR), i.e.:
@@ 3 What is the real “Hardness” of CQA in incomplete KGs?
Given a query $q$ and every answer set of candidate answers $t\in\mathcal{E}_{q}$, we define the reasoning tree of each QA pair $(q,t)$, as the directed acyclic graph starting from the anchor entities of $q$ to the target entity $t$, whose relational structure matches the query graph. Figure 1 provides examples of different reasoning trees for four different answers to the same query. There, we highlight whether a link belongs to $G_{\text{train}}$ or not, i.e., it is a missing link. We assess the hardness of each answer $t\in\mathcal{E}_{q}$, by analyzing the composition of the reasoning tree required to predict $t$. As to answering a $(q,t)$ pair, multiple reasoning trees are possible, so we select the one with the smallest number of missing links and the fewest number of hops. A $(q,t)$ pair is trivial if the answer $t$ can be entirely retrieved from $G_{\text{train}}$, i.e., there exists at least one reasoning tree where all the links in it are seen during training. This type of answer does not need probabilistic inference and hence is filtered out from current CQA benchmarks (Ren & Leskovec, 2020), which only consider non-trivial $(q,t)$ pairs, which we call inference-based pairs. However, inference-based pairs do not need all links in their reasoning tree to be predicted as, by definition, it is sufficient that at least one link in the tree is missing. Therefore, we define a $(q,t)$ pair to be a partial-inference pair if its reasoning tree contains at least one link in $G_{\text{train}}$ and at least one link present in $G_{\text{test}}$. Alternatively, a $(q,t)$ pair whose reasoning tree contains only links present in $G_{\text{test}}$ is called a full-inference pair. We claim that inference-based queries encompass a spectrum of hardness, with full-inference queries being the hardest and queries that have only one truly missing link (1p) the easiest. However, current benchmarks flatten this complexity and highlight only performance on the easiest.
@@ 4 How Many “Easy” Queries in Current CQA Benchmarks?
In this section, we systematically analyze the practical hardness of queries from the current CQA benchmarks and answer the following research question: (RQ1) What is the proportion of QA pairs that can be classified as full-inference rather than partial-inference and how easy are the latter? To this end, we consider the CQA benchmarks generated from $\mathsf{FB15k237}$ (based on Freebase) and $\mathsf{NELL995}$ (based on NELL systems (Carlson et al., 2010)) as they are the most used to evaluate SoTA methods for CQA (Ren et al., 2020; Ren & Leskovec, 2020; Arakelyan et al., 2021; Zhang et al., 2021; Zhu et al., 2022; Arakelyan et al., 2023). [^4] We do not consider $\mathsf{FB15k}$ (Bordes et al., 2013) as it suffers from another form of data leakage (Toutanova & Chen, 2015).
Complex queries can be reduced to much simpler types. We group the testing QA pairs into query types (Sec. 2), and we further split them based on whether they can be reduced to simpler types after observing the training links in $G_{\operatorname{train}}$ (Sec. 3). Table 1 shows that for both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ the vast majority of QA pairs can be reduced to simpler types. For $\mathsf{FB15k237}$, 86.8% to 98.3% of QA pairs can be reduced to 1p queries, while only 0.1% to 4% require full inference. Similarly, for $\mathsf{NELL995}$, 49.5% to 98.5% of QA pairs map to 1p queries, and only 0.1% to 6% to full inference. For instance, 96.7% of 2i1p QA pairs in $\mathsf{FB15k237}$ can be reduced to 1p queries, 1.8% to 2p, and 1.4% to 2i. However, only 0.1% of these pairs cannot be reduced to any other QA pair, i.e., they require full inference in order to be predicted. The only exceptions to this trend are QA pairs where the query has a 1p or 2u structure which, by definition, only require the prediction of a single link and therefore cannot be reduced by any other query type. We conduct the same analysis for queries that include negation, considering only the reasoning tree over non-negated links (see Sec. A.1 for details). Table A.1 presents results analogous to those in Table 1, showing that the vast majority of QA pairs including negation have training links in their reasoning tree. For example, 95.4% and 93.9% of 3in QA pairs in $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, respectively, have training links in their reasoning tree.
@@
Non-existing links for union queries. Finally, we discovered that there are non-existing links, i.e., links that are not in the original KG $G$ and hence neither in $G_{\operatorname{train}}$ or $G_{\operatorname{test}}$, in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ in the reasoning trees of queries involving unions, i.e., the 2u and 2u1p types. These links violate our definitions of inference QA pairs, hence, we filter them out, and report only filtered QA pairs in Table 1. For example, for 2u QA pairs, if either one of the two links in the reasoning tree does not exist in the graph, we remove the QA pair (see Figure 3 bottom for an example). It follows that only QA pairs where both links in the reasoning tree exist in the test graph are retained (see Figure 3 (Top) for an example). More crucially, these non-existing links can alter the performance of solvers for 2u and 2u1p types, as we discuss next.
Takeaway 1. We discourage using current CQA benchmarks, as they essentially reflect only the performance of predicting answers where only one link is truly missing (for both hybrid and neural solvers).
@@
CLMPT 11.3 11.3 6.1 4.0 - - - - 19.3 19.1 18.7 - 7.9 - - 13.2
Table 2: SoTA performance drops significantly on full-inference QA pairs as shown for the MRR for 3p and 2i1p queries on $\mathsf{FB15k237}$. Compare the overall MRR achieved on all available testing queries (column “ovr”) with that achieved on the 1p reductions (“1p” column, i.e., when only a single link is truly missing): they are extremely close. This is because
Figure 3: Some answers of union queries can be reached by a single missing link, while the other link does not exist, as shown for this 2u query and fragments of the KG $\mathsf{FB15k237}$, where $r_{1}=\mathsf{releasedIn}$, $r_{2}=\mathsf{crewMemberOf}$. (Top) Example of an answer that we retain in our benchmark, being reachable by two missing links. (Bottom) Example of answers we filter out, being only reachable by one missing link.
@@ 5 What do SoTA Models Perceive as Hard?
Neural CQA methods. Over the years, a significant number of neural models have been proposed for solving the CQA task, see App. C for an overview. Among purely-neural models, we consider five representative approaches that significantly differ in their methodological designs and yield SoTA results compared to other models in their class: (1) Cone Embeddings (ConE; Zhang et al., 2021) is a geometry-based complex query answering model where logical conjunctions and disjunctions are represented as intersections and union of cones. (2) Graph Neural Network Query Executor (GNN-QE; Zhu et al., 2022) decomposes a complex first-order logic query into projections over fuzzy sets. (3) Conditional Logical Message Passing Transformer (CLMPT; Zhang et al., 2024) leverages existing KG embeddings to conduct one-hop inferences on atomic formulas in a message-passing scheme as in LMPNN (Wang et al., 2023), and extends it by using a transformer to aggregate messages incoming from neighbor nodes. (4) Continuous Query Decomposition (CQD; Arakelyan et al., 2021, 2023), reduces the CQA task to the problem of finding the most likely variable assignment, where the likelihood of each link (1p query) is assessed by a neural link predictor, and logical connectives are relaxed via fuzzy logic operators. (5) UltraQuery (ULTRAQ; Galkin et al., 2024b) is a foundation model for CQA inspired by Galkin et al. (2024a) where links and logical operations are represented by vocabulary-independent functions which can generalize to new entities and relation types in any KG.
Hybrid solvers. So far, we have assessed that current benchmarks are skewed towards easier query types (Table 1) and thus the perceived progress of current SoTA CQA methods boils down to their performance on 1p queries (Table 2). To have an undistorted view of this progress, in the next section, we will devise a benchmark that rebalances all partial-inference as well as full-inference queries types. That is, we will provide an equal number of QA pairs for each sub-type a query can be reduced to. Now, we argue that in a real-world scenario one has to perform reasoning over both existing links and missing ones, hence discarding the information in $G_{\operatorname{train}}$ is wasteful. This, in addition to the fact that purely-neural SoTA CQA methods might implicitly exploit this aspect at test time if they are able to memorize well $G_{\operatorname{train}}$, pushes us to evaluate also hybrid solvers.
Table 2 presents a portion of the results in terms of MRR for the testing QA pairs in the existing benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, along with the different QA pairs they can be reduced to when observing the training KG, grouped by each query type. There, the performance of all SoTA methods consistently drops for query answer pairs that have a higher number of missing links in their reasoning tree. Furthermore, Tables A.4 and A.5 compare the MRR computed on all the available queries originally used in the benchmarks (column “all”), and the scores on those query types that they can be reduced to. There is a high similarity in MRR between the columns “all” and “1p” that is evidence that the good performances of CQA methods are explained by their link prediction performances, as a very high percentage of queries in fact are reduced to 1p queries (see Table 1). A similar conclusion can be drawn for queries including negation, as shown in Table A.2.
Note that this happens consistently also for hybrid solvers such as QTO and our CQD-Hybrid: they have a harder time solving QA pairs with more than one missing link. While their performance was consistently surpassing purely-neural solvers, and CQD-Hybrid consistently boosted the performance of CQD, their poor performance on harder QA pairs is not reflected in the overall average score that is usually reported on papers, see Table 2.
Table 3: Exploiting training links during inference boosts MRR on existing benchmarks, as shown by our CQD-Hybrid model compared to previous SoTA, for several query types. CQD-Hybrid always achieves higher MRR scores when compared to CQD, and outperforms more sophisticated non-hybrid baselines 9/14 times. Best values in bold and second best underlined.
Model 2p 3p 2i 3i 1p2i 2i1p 2u1p
$\mathsf{FB15k237}$ GNN-QE 14.7 11.8 38.3 54.1 31.1 18.9 9.7
ULTRAQ 11.5 8.9 35.7 51.0 29.6 18.6 7.3
CQD 13.2 7.8 35.0 48.5 27.5 20.7 10.5
ConE 12.8 11.0 32.6 47.3 25.5 14.0 7.4
CLMPT 13.9 11.3 37.5 51.9 28.4 19.3 11.4
CQD-Hybrid 15.0 11.0 37.6 52.7 31.2 24.0 10.3
$\mathsf{NELL995}$ GNN-QE 17.9 15.2 40.0 50.9 29.1 20.5 8.8
ULTRAQ 11.2 9.7 36.3 47.7 25.1 15.6 8.4
CQD 22.0 13.4 42.2 51.8 31.5 25.8 16.0
ConE 16.0 13.8 39.6 50.2 26.1 17.6 11.1
CLMPT 22.1 18.2 41.6 51.7 29.0 24.7 16.1
CQD-Hybrid 23.8 17.8 44.2 57.8 33.2 28.4 16.6
Table 3 highlights that our CQD-hybrid is able to surpass existing non-hybrid SoTA, which is evidence that a pre-trained neural link predition plus memorization of $G_{\operatorname{train}}$ achieves SoTA performances for $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. The complete stratified comparison for all solvers is reported in Tables A.4 and A.5. However, in a few cases it happens that full-inference results are higher than some partial-inference; in such cases, the reason is to be found in the fact that full-inference QA pairs are very scarce (see Table 1). [^5] We rule out that this is due to the influence of the number of existing entities bounding to existential quantified variables. For example we report such an analysis for CQD in Figures A.2 and D.2. Moreover, we found that no anchor entities nor relation names is predominant in both benchmarks, as shown in Table A.6. This motivates us to create a fully-balanced benchmark in the next section.
@@
CLMPT 45.3 14.8 32.8 58.1 18.6 39.6
@@
Non-existing links for union queries give a false sense of hardness. For union queries, we filter query answer pairs, as discussed in Sec. 4, removing those with non-existing links in their reasoning trees. In Table 4, we show that if we do not do that, and consider also the pairs with non-existing links (denoted as “2u all”), MRR performance greatly drops. In this way, we reproduce the low performance for 2u queries that the original SoTA baselines reported in their papers. However, as Table 4 shows, this is just an artifact due to including non-filtered pairs while computing Equation 1. With our filtered pairs, instead, we recover the performance of 1p queries, as expected (Sec. 2). Similar conclusions can be drawn for other union queries as reported in Table A.3.
@@
Takeaway 2. $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess CQA methods, as most of their QA pairs can be predicted by just predicting a single missing links. This results in highly inflated performance that distorts the perception of progress in the field and pushes the community to improve the performance only on the easiest QA pairs.